Hello! This is all one question, thank you very much in advance!

AFN EQUATION

Carlsbad Corporation's sales are expected to increase from $5 million in 2016 to $6 million in 2017, or by 20%. Its assets totaled $3 million at the end of 2016. Carlsbad is at full capacity, so its assets must grow in proportion to projected sales. At the end of 2016, current liabilities are $1 million, consisting of $250,000 of accounts payable, $500,000 of notes payable, and $250,000 of accrued liabilities. Its profit margin is forecasted to be 3%.

  1. Assume that the company pays no dividends.
    Under these assumptions, what would be the additional funds needed for the coming year? Write out your answer completely. For example, 5 million should be entered as 5,000,000. Round your answer to the nearest cent.
    $

  2. Why is this AFN different from the one when the company pays dividends?
    1. Under this scenario the company would have a lower level of retained earnings, which would decrease the amount of additional funds needed.
    2. Under this scenario the company would have a higher level of retained earnings, which would reduce the amount of additional funds needed.
    3. Under this scenario the company would have a higher level of retained earnings, which would reduce the amount of assets needed.
    4. Under this scenario the company would have a higher level of spontaneous liabilities, which would reduce the amount of additional funds needed.
    5. Under this scenario the company would have a lower level of retained earnings, which would increase the amount of additional funds needed.
Answer #1

1: AFN = Increase in assets-Increase in spontaneous liabilities -Retained earnings

Increase in assets =3000000*20%= 600000

Increase in spontaneous liabilities= (250000+250000)*20%=100000

Retained earnings = Profit margin*Sales- Dividend = 3%*6000000=180000

AFN = 600000-100000-180000 = 320000

2: Option 5

When the company pays dividends, the retained earnings are lesser and hence the amount of additional funds required will be more since internal funds will be lesser.

Option 4 is incorrect because paying dividends will not change the spontaneous liabilities. Option 1 is incorrect because paying dividends will decrease the retained earnings due to which external funds will be required more. Option 2 and 3 are incorrect because paying dividends will not increase retained earnings rather decrease them.